The separation between a node and the next antinode in a vibration air column is 25 cm. If the speed of sound in air is 340 m s-1, find the frequency of vibration of the air column.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. Here given I base 2 = 0.67 m, l base 1 = 0.2 m, f = 400 Hz We know that λ = 2(l base 2 – l base 1) ⇒ λ = 2(62 – 20) = 84 cm = 0.84 m. So, v = nλ = 0.84 × 400 = 336 m/s We know from above that, L base 1 + d = λ/4 ⇒ d = λ/4 – l base 1 = 21 – 20 = 1 cm.